OpenStudy (zab505):

A colony of bacteria is growing exponentially according to the function below, where t is in hours. What will the approximate number of bacteria be after 8 hours? B(t)=4*e^0.8t A. 2407 B. 205,968 C. 602 D. 4,289,241

OpenStudy (aaronq):

Plug t=8 h into the formula \(\sf B(t)=4*e^{0.8t}\) \(\sf B(8)=4*e^{0.8*8}=4525.1816935371888104\) you might want to round it

OpenStudy (aaronq):

wait that doesnt match any of the answers. hold on

OpenStudy (aaronq):

i made a mistake when i used the calculator, it should be: \(\sf B(8)=2407.380\)

OpenStudy (zab505):

Wow thank you so much!

OpenStudy (aaronq):

no problem man

OpenStudy (zab505):

Hey do you think you can help on this one?

OpenStudy (aaronq):

I think the have the same horizontal translation, both have a -2 x-translation.

OpenStudy (zab505):

Thank you!

OpenStudy (aaronq):

no probs!